HUBBLE SPACE TELESCOPE DAILY REPORT #1765 PERIOD COVERED: 0000Z (UTC) 11/21/96 - 0000Z (UTC) 11/22/96 Daily Status Report as of 327/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.2 Completed FOS/BL 6683 (Eclipsing Binaries in the Magellanic Clouds = - Fundamental Properties A) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of O/B Class, eclipsing binary system star = SMC-HV1620 located in the Small Magellanic Cloud. From these data and = the latest Kurucz model atmospheres, accurate temperatures will be = derived. The observations were executed as scheduled, and no anomalies = were noted. 1.3 Completed FOS/BL 6799 (The BAL Region Covering Factor in a Sample = of IRAS-Selected QSOs) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of QSO MARK231 in order to determine the = criteria by which the presence of BALs in the spectra may be predicted. = The observations were completed as planned, and no problems were = reported. 1.4 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.5 Completed FGS 6921 (Intermediate Plate Scale Calibration) The FGSs were used in the POS Mode to make astrometric = observations of star GRW+77D582 and several other reference stars. These = data will be used for an intermediate plate scale calibration of the = central region of the FGSs, where all the annual parallax work is done. = The reference stars were taken from the HIPPARCOS observing list. The = observations were executed as scheduled, and no anomalies were noted. 1.6 Completed FOS/RD 6925 (FOS Observations of Solar Analog Standards) The Faint Object Spectrograph (Red Side) was used to make absolute = spectrophotmetry observations of the solar analog star P041C-cali in the = spectral range of 2200 A0 to 8500 A0. The objective is to provide the = primary standards for the absolute calibration of the NICMOS. The = observations were completed as scheduled, and no problems were reported. 1.7 Completed WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with = the prime science observations as part of the Medium Deep Sky Survey. = The observations were executed as scheduled, and no problems were = reported. 1.8 Completed FOS/BL 6546 (Dissecting the wind and Disk of the = Nova-like Variable V347 Puppus) The Faint Object Spectrograph (Blue Side) was used to make = observations of the nova-like, variable star V347 Puppus in order to = study its disk and to dissect its winds. The ultraviolet resonance = spectral lines indicate that this type of cataclysmic variable suffers a = high mass loss rate in the form of high velocity winds. The observations = were completed as scheduled, and no problems were reported. 1.9 Completed Two Sets of WF/PC-2 7203 (Cosmology with the Medium Deep = Survey - Short Orbit Cases) The WF/PC-2 was used to take short (about 20 minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems = were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: There were no significant operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH